OpenStudy (anonymous):

Name the property of real numbers illustrated by the equation. 5+(sqrt 15+6)=(5+sqrt15)+6

OpenStudy (fibonaccichick666):

an inaccurate one

OpenStudy (anonymous):

Associative Property of Addition Commutative Property of Multiplication Associative Property of Multiplication

OpenStudy (fibonaccichick666):

\[5+\sqrt (15+6)=(5+\sqrt(15) )+6\] right?

OpenStudy (fibonaccichick666):

or is it? \[5+(\sqrt (15)+6)=(5+\sqrt(15) )+6\]

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

which one?

OpenStudy (anonymous):

second

OpenStudy (anonymous):

\[5+(\sqrt{15}+6)=(5+\sqrt{15})+6\]

OpenStudy (anonymous):

its commutative property of multiplication right?

OpenStudy (fibonaccichick666):

no are we multiplying anything?

OpenStudy (anonymous):

no

OpenStudy (fibonaccichick666):

ok so if we are not multiplying what are we doing?

OpenStudy (anonymous):

adding

OpenStudy (fibonaccichick666):

right so we have two properties for addition that we could be using(depending on the definition), can you tell me their names?

OpenStudy (anonymous):

idk

OpenStudy (fibonaccichick666):

what are the properties?

OpenStudy (fibonaccichick666):

in general for addition